# The Fast and the Frame-Dragging: Efficient waveforms for asymmetric-mass eccentric equatorial inspirals into rapidly-spinning black holes

### v1

@@ Appendix E Comparison against other implementations in limiting cases
In addition to the self-contained examination of our few v2 carried out in Section IV, we can also verify that it correctly matches other models in limiting cases. Specifically, we compare against accurate models for quasi-circular inspirals into spinning MBHs Khalvati:2024tzz ; Nasipak:2023kuf (Section E.1), and verify that elements of our model approach PN-GSF based results in the weak-field limit ($p\gg p_{\mathrm{sep}}$) with moderate eccentricities (Section E.2).
@@ E.1 Asymmetric-mass, quasi-circular inspirals into rapidly-spinning black holes
Before computing the phase shifts, we rescaled the masses as discussed in Appendix A to have a similar mass convention across all three models. Moreover, since the overall scaling of the phases, and phase errors are known with the mass ratio and with the primary mass, we only report the comparison as a function of the primary spin. We chose masses of $(m_{1},m_{2})=(10^{6},10)M_{\odot}$. Throughout this section, for each value of the primary spin parameter $a$, the inspiral is started at a corresponding $p_{0}$ such that the evolution lasts four years before reaching the plunge.
@@
The absolute value of the final phase shift as a function of primary spin compared to both models is shown in Fig. 22. The phase shift $\Delta\Phi_{\phi}\lesssim 1$ for all spin values; we observe some oscillations alongside the overall increase for higher spins in prograde orbits. The primary source of these phase shifts lies in differences in flux and interpolation accuracy. These discrepancies arise from several factors, such as the accuracy of the flux data, choice of parameterization, the scaling applied to the flux data prior to interpolation, the type of interpolant used, and the interpolation grid itself, specifically, the density of said grid.
@@
The KerrCirc model employs a uniform grid in spin with $\Delta a=0.01$, whereas BHPWave uses a non-uniform grid with a higher density of points concentrated at larger spin values. The latter setup is more similar to ours, in which the grid resolution is denser at higher spin and sparser at lower spin. This grid structure likely explains why BHPWave shows better agreement with our model. A detailed study of grid-induced and interpolation-induced errors as a function of spin is under investigation khalvati2025sys .
@@
Figure 23: Comparison of the real part of the mode amplitude differences, $|\Delta\mathcal{A}_{\ell m0}|$, as a function of orbital separation $p$. Each panel corresponds to a fixed primary spin value, $a\in{0.99,0.5,-0.5,-0.99}$, and shows three different modes, $(\ell,m)\in{(2,2),(5,5),(10,10)}$. The solid lines compare our model to KerrCirc, while the dashed lines compare KerrCirc to BHPWave. The innermost stable circular orbit is indicated by a vertical marker.
@@
Furthermore, we compared the interpolated amplitudes in our model for four different primary spin values, $a\in\{0.99,0.5,-0.5,-0.99\}$ and three different modes, $(\ell,m)\in\{(2,2),(5,5),(10,10)\}$ as a function of orbital separation, $p$, with both KerrCirc, and BHPWave in Fig. 23. We dropped the index $n$ here because in the circular orbit limit, there are no radial harmonics and $n=0$. We plot $|\Delta\mathcal{A}_{\ell m0}|$ instead of the fractional error, as the latter can appear misleadingly large for higher modes due to their smaller amplitudes. We observe that the mode amplitudes agree more closely between KerrCirc and BHPWave than with our model. This is primarily because we do not employ a very dense grid in $p$ for computing the amplitudes (Appendix B). We remind the reader that amplitude errors do not accumulate over the inspiral and therefore do not lead to significant deviations in the total mismatch between waveforms (this is observed during amplitude validation in Section IV.2).
@@
Finally, to provide an overall comparison of the full waveforms, we computed the mismatch (see Appendix D for definitions) between our waveform and both KerrCirc and BHPWave as a function of the primary spin for $a\in[-0.99,0.99]$ in the circular orbit limit. The mismatches were calculated using a flat PSD, such that differences between models are captured uniformly across all frequencies. As for the phase shifts, we fixed the source masses to $(m_{1},m_{2})=(10^{6},10)M_{\odot}$ and viewing angles are fixed to typical value for the viewing angles $(\theta,\phi)=(\pi/5,\pi/3)$. The results are shown in Fig. 24. The mismatch between our waveform and BHPWave remains below $10^{-5}$ for almost all spin values, in both prograde and retrograde orbits, except for a sharp increase as $a\rightarrow 0.99$. This increase coincides with the region where the phase shifts and amplitude differences are largest, as previously discussed. Typically, for larger spin values, as the system enters deeper into the strong-field regime, any discrepancies sourced from differences in amplitudes, fluxes, grid structures, or interpolation schemes become more pronounced.
@@
In comparison, the mismatch with KerrCirc is systematically larger, which is consistent with the phase shift trends observed in Fig. 22. It is worth mentioning that these mismatch values do not necessarily reflect the absolute accuracy of any of the models considered here, but rather indicate the level of agreement between them. In most cases, the discrepancies are small enough to be below the resolvable threshold of the detector and thus unlikely to impact observational outcomes. A detailed investigation into the sources of these differences is beyond the scope of this article and will be presented in a forthcoming study focusing on circular equatorial waveforms khalvati2025sys .
@@ E.2 Weak-field behaviour against PN-GSF approaches
The combined PN-GSF approach of black hole perturbation theory yields closed-form analytic solutions to the (frequency-domain) Teukolsky equation and thus the inspiral forcing functions in Eq. 2 and the GW strain amplitudes in Eq. 5 in turn; we refer a reader to Refs. Mino:1997bx ; Sasaki:2003xr for reviews. In the context of the EMRI and IMRI modelling, this approach is dependent upon order-by-order expansions in powers of the (dimensionless) velocity parameter: $v=\sqrt{1/p}$ (or similar parameters) and often eccentricity: $e$ (while keeping the inclination and primary spin arbitrary), limiting the accuracy of the analytic expressions in the stronger field ($v\simeq 1$) and at larger eccentricities ($e\simeq 1$). At the same time, this method is particularly advantageous to account for a vast region of the weak-field parameter space ($v\ll 1$) that would be expensive to cover with numerical GSF computations.
@@
Building upon the results reported in Sago:2005fn ; Ganz:2007rf ; Sago:2015rpa , Ref. Fujita:2020zxe derived the analytic expressions for the inspiral forcing functions equivalent to Eq. 2 for generic (eccentric and precessing) adiabatic inspirals, valid to $\mathcal{O}(v^{11},e^{11})$ relative to the leading “Newtonian-circular” terms. Shortly afterwards, Ref. Isoyama:2021jjd combined that result with the analytic GW strain amplitudes (derived to the same PN accuracy) to produce a generic adiabatic waveform model. In this work, we label these analytical results simply “PN5” for brevity but we highlight the additional eccentricity expansion employed in their derivation. The PN5 expressions for both the spheroidal harmonic modes of the Teukolsky amplitudes and the inspiral forcing functions are publicly available from the Black Hole Perturbation Club (BHPC) website BHPClub  [^11] There are a number of other PN-GSF calculations that explicitly consider the eccentric equatorial inspirals in Kerr Tagoshi:1995sh ; Bini:2016dvs ; Bini:2019lcd ; Munna:2023wce . However, the results obtained are lower orders in $p$ and $e$ expansion (sometimes with slow primary spin $a\ll 1$ assumed), or have yet to be implemented in few. For this reason, we have not used them as a weak-field benchmark for the results in this work..
@@
Figure 25: The relative error in $\hat{f}^{(0)}_{p,e}$ between the interpolated forcing function in this work (few, $\hat{f}^{\mathrm{FEW}}_{p,e}$) and the PN forcing function in Section E.2 (PN5, $\hat{f}^{\mathrm{PN5}}_{p,e}$) in the weak-field for $a=0.998$ (we find similar behaviour for other values of $a$).
@@
Figure 26: Absolute relative error for three selected modes of the interpolated (complex) mode amplitude in this work (few, ${\mathcal{A}}^{\mathrm{FEW}}_{\ell mn}$) and the PN5 waveform mode amplitude in Section E.2 (PN5, ${\mathcal{A}}^{\mathrm{PN5}}_{\ell mn}$) with $a=0.998$ (prograde); note the difference in the range of $p$ in the right-bottom panel, which depicts the error in the strong field region. We find similar levels of agreement across other values of $a$.
@@
Figure 27: (From left to right): Orbital dephasings between the trajectory model presented in this work (Kerr) and the PN trajectory discussed in Section E.2 (PN5) for $m_{1}\in\{10^{5},10^{6},10^{7}\}\,M_{\odot}$. All inspirals are evolved for $4$ years, with $a=0.998$ and $\epsilon=10^{-5}$. The white dashed contours indicate regions where the dephasing crosses $\{0.01,0.1,1\}\,\mathrm{rad}$. The blue dashed line indicates the outer boundary in $(p_{0},e_{0})$ where the body plunges within four years in the PN5 model.
@@
Some of these PN5 results have already been integrated into few. For instance, the PN5-AAK waveform model combines the PN5 inspiral trajectory with the semi-relativisitic (“kludge”) amplitudes chua2017augmented as described in Ref. Katz:2021yft , providing a semi-relativistic waveform model defined over the generic inspiral parameter space: the few implementation of the PN5 mode amplitudes is also under development (cf. Section VI.2).
@@
    Validating the interpolations of the forcing functions $\hat{f}^{(0)}_{p,e}$ in the trajectory module against their PN5 equivalents (Fig. 25).
@@
    Validating the interpolations of the mode amplitudes $\mathcal{A}_{\ell mn}$ against their PN5 equivalents for sample harmonic modes; the PN5 spherical harmonic waveform mode amplitudes were obtained from the PN5 Teukolsky spheroidal harmonic mode amplitudes via Eq. 6 and projected onto spherical harmonics as described in Eq. 11 (Fig. 26).
@@
    Assessing the orbital dephasing (over four years) between the (fully-relativistic) trajectory model in few v2 and the PN5 inspiral model scales(Fig. 27).
To do this, we have since implemented two minor improvements to the PN5 trajectory module. Firstly, we have corrected minor typos in the (adiabatic) PN5 fluxes of the angular momentum and Carter constant and the associated PN5 forcing functions $\hat{f}^{\mathrm{PN5}}_{p,e}$. Secondly, unlike the public database at BHPC BHPClub , $\hat{f}^{\mathrm{PN5}}_{p,e}$ here are computed from the PN5 fluxes of energy and angular momentum (in the equatorial limit) without expanding the Jacobian elements in Eq. 7 in powers of $v$ or $e$, as doing so enhances the accuracy of $\hat{f}^{\mathrm{PN5}}_{p,e}$ in comparison to their numerical counterparts (cf. Ref. Isoyama:2021jjd ).
We find in Fig. 25 that the agreement between the interpolations of the forcing functions $\hat{f}^{\mathrm{FEW}}_{p,e}$ and the PN5 forcing functions $\hat{f}^{\mathrm{PN5}}_{p,e}$ improves as $p$ increases and $e$ decreases. This agreement in ${\hat{f}}^{(0)}_{e}$ breaks down for small $e$, since both magnitudes of $\hat{f}^{\mathrm{PN5,Few}}_{e}$ become very small. A similar trend can be observed in the amplitude comparisons displayed in the top two panels of Fig. 26, aside from the bottom left panel where the magnitude of the higher multipole amplitude varies rapidly and become small, and the bottom right panel where the accuracy of the PN5 amplitudes decrease closer to the separatrix. However, we find the interpolation error is predominant at smaller values of $e$ in this case. The deterioration of the agreement in ${\mathcal{A}}_{\ell mn}$ as the eccentricity becomes larger (typically $e\gtrsim 0.4$) is a result of the eccentricity expansion in the PN5 amplitudes, which become significantly less accurate in that region. This observation has been corroborated through a comparison of the PN5 amplitudes with numerical amplitude data from both the BHPT BHPToolkit and the BHPC Isoyama:2021jjd ; BHPClub . For weak-field trajectories with initial separation $p_{0}\gg p_{\text{sep}}$ in Fig. 27, we observe that the dephasing between both models tends to zero (independent of the initial eccentricity $e_{0}$), indicating that the two models converge in the weak field limit, as expected.

### v2

@@ Appendix E Comparison against other implementations in limiting cases
In addition to the self-contained examination of our few v2 carried out in Section˜IV, we can also verify that it correctly matches other models in limiting cases. Specifically, we compare against accurate models for quasi-circular inspirals into spinning MBHs Khalvati:2024tzz ; Nasipak:2023kuf (Section˜E.1), and verify that elements of our model approach PN-GSF based results in the weak-field limit ($p\gg p_{\mathrm{sep}}$) with moderate eccentricities (Section˜E.2).
@@ E.1 Asymmetric-mass, quasi-circular inspirals into rapidly-spinning black holes
Before computing the phase shifts, we rescaled the masses as discussed in Appendix˜A to have a similar mass convention across all three models. Moreover, since the overall scaling of the phases, and phase errors are known with the mass ratio and with the primary mass, we only report the comparison as a function of the primary spin. We chose masses of $(m_{1},m_{2})=(10^{6},10)M_{\odot}$. Throughout this section, for each value of the primary spin parameter $a$, the inspiral is started at a corresponding $p_{0}$ such that the evolution lasts four years before reaching the plunge.
@@
The absolute value of the final phase shift as a function of primary spin compared to both models is shown in Fig.˜22. The phase shift $\Delta\Phi_{\phi}\lesssim 1$ for all spin values; we observe some oscillations alongside the overall increase for higher spins in prograde orbits. The primary source of these phase shifts lies in differences in flux and interpolation accuracy. These discrepancies arise from several factors, such as the accuracy of the flux data, choice of parameterization, the scaling applied to the flux data prior to interpolation, the type of interpolant used, and the interpolation grid itself, specifically, the density of said grid.
@@
The KerrCirc model employs a uniform grid in spin with $\Delta a=0.01$, whereas BHPWave uses a non-uniform grid with a higher density of points concentrated at larger spin values. The latter setup is more similar to ours, in which the grid resolution is denser at higher spin and sparser at lower spin. This grid structure likely explains why BHPWave shows better agreement with our model. A detailed study of grid-induced and interpolation-induced errors as a function of spin is reported in Khalvati:2025znb .
@@
Figure 23: Comparison of the real part of the mode amplitude differences, $|\Delta\mathcal{A}_{\ell m0}|$, as a function of orbital separation pp. Each panel corresponds to a fixed primary spin value, $a\in{0.99,0.5,-0.5,-0.99}$, and shows three different modes, $(\ell,m)\in{(2,2),(5,5),(10,10)}$. The solid lines compare our model to KerrCirc, while the dashed lines compare KerrCirc to BHPWave. The innermost stable circular orbit is indicated by a vertical marker.
@@
Furthermore, we compared the interpolated amplitudes in our model for four different primary spin values, $a\in\{0.99,0.5,-0.5,-0.99\}$ and three different modes, $(\ell,m)\in\{(2,2),(5,5),(10,10)\}$ as a function of orbital separation, $p$, with both KerrCirc, and BHPWave in Fig.˜23. We dropped the index $n$ here because in the circular orbit limit, there are no radial harmonics and $n=0$. We plot $|\Delta\mathcal{A}_{\ell m0}|$ instead of the fractional error, as the latter can appear misleadingly large for higher modes due to their smaller amplitudes. We observe that the mode amplitudes agree more closely between KerrCirc and BHPWave than with our model. This is primarily because we do not employ a very dense grid in $p$ for computing the amplitudes (Appendix˜B). We remind the reader that amplitude errors do not accumulate over the inspiral and therefore do not lead to significant deviations in the total mismatch between waveforms (this is observed during amplitude validation in Section˜IV.2).
@@
Finally, to provide an overall comparison of the full waveforms, we computed the mismatch (see Appendix˜D for definitions) between our waveform and both KerrCirc and BHPWave as a function of the primary spin for $a\in[-0.99,0.99]$ in the circular orbit limit. The mismatches were calculated using a flat PSD, such that differences between models are captured uniformly across all frequencies. As for the phase shifts, we fixed the source masses to $(m_{1},m_{2})=(10^{6},10)M_{\odot}$ and viewing angles are fixed to typical value for the viewing angles $(\theta,\phi)=(\pi/5,\pi/3)$. The results are shown in Fig.˜24. The mismatch between our waveform and BHPWave remains below $10^{-5}$ for almost all spin values, in both prograde and retrograde orbits, except for a sharp increase as $a\rightarrow 0.99$. This increase coincides with the region where the phase shifts and amplitude differences are largest, as previously discussed. Typically, for larger spin values, as the system enters deeper into the strong-field regime, any discrepancies sourced from differences in amplitudes, fluxes, grid structures, or interpolation schemes become more pronounced.
@@
In comparison, the mismatch with KerrCirc is systematically larger, which is consistent with the phase shift trends observed in Fig.˜22. It is worth mentioning that these mismatch values do not necessarily reflect the absolute accuracy of any of the models considered here, but rather indicate the level of agreement between them. In most cases, the discrepancies are small enough to be below the resolvable threshold of the detector and thus unlikely to impact observational outcomes. A detailed investigation into the sources of these differences is beyond the scope of this article and is presented separately in a study focusing on the systematics errors in circular equatorial waveforms Khalvati:2025znb .
@@ E.2 Weak-field behaviour against PN-GSF approaches
The combined PN-GSF approach of black hole perturbation theory yields closed-form analytic solutions to the (frequency-domain) Teukolsky equation and thus the inspiral forcing functions in Eq.˜2 and the GW strain amplitudes in Eq.˜5 in turn; we refer a reader to Refs. Mino:1997bx ; Sasaki:2003xr for reviews. In the context of the EMRI and IMRI modelling, this approach is dependent upon order-by-order expansions in powers of the (dimensionless) velocity parameter: $v=\sqrt{1/p}$ (or similar parameters) and often eccentricity: $e$ (while keeping the inclination and primary spin arbitrary), limiting the accuracy of the analytic expressions in the stronger field ($v\simeq 1$) and at larger eccentricities ($e\simeq 1$). At the same time, this method is particularly advantageous to account for a vast region of the weak-field parameter space ($v\ll 1$) that would be expensive to cover with numerical GSF computations.
@@
Building upon the results reported in Sago:2005fn ; Ganz:2007rf ; Sago:2015rpa , Ref. Fujita:2020zxe derived the analytic expressions for the inspiral forcing functions equivalent to Eq.˜2 for generic (eccentric and precessing) adiabatic inspirals, valid to $\mathcal{O}(v^{11},e^{11})$ relative to the leading “Newtonian-circular” terms. Shortly afterwards, Ref. Isoyama:2021jjd combined that result with the analytic GW strain amplitudes (derived to the same PN accuracy) to produce a generic adiabatic waveform model. In this work, we label these analytical results simply “PN5” for brevity but we highlight the additional eccentricity expansion employed in their derivation. The PN5 expressions for both the spheroidal harmonic modes of the Teukolsky amplitudes and the inspiral forcing functions are publicly available from the Black Hole Perturbation Club (BHPC) website BHPClub  [^11] There are a number of other PN-GSF calculations that explicitly consider the eccentric equatorial inspirals in Kerr Tagoshi:1995sh ; Bini:2016dvs ; Bini:2019lcd ; Munna:2023wce . However, the results obtained are lower orders in $p$ and $e$ expansion (sometimes with slow primary spin $a\ll 1$ assumed), or have yet to be implemented in few. For this reason, we have not used them as a weak-field benchmark for the results in this work..
@@
Figure 25: The relative error in $\hat{f}^{(0)}_{p,e}$ between the interpolated forcing function in this work (few, $\hat{f}^{\mathrm{FEW}}_{p,e}$) and the PN forcing function in Section˜E.2 (PN5, $\hat{f}^{\mathrm{PN5}}_{p,e}$) in the weak-field for $a=0.998$ (we find similar behaviour for other values of $a$).
@@
Figure 26: Absolute relative error for three selected modes of the interpolated (complex) mode amplitude in this work (few, ${\mathcal{A}}^{\mathrm{FEW}}_{\ell mn}$) and the PN5 waveform mode amplitude in Section˜E.2 (PN5, ${\mathcal{A}}^{\mathrm{PN5}}_{\ell mn}$) with $a=0.998$ (prograde); note the difference in the range of $p$ in the right-bottom panel, which depicts the error in the strong field region. We find similar levels of agreement across other values of $a$.
@@
Figure 27: (From left to right): Orbital dephasings between the trajectory model presented in this work (Kerr) and the PN trajectory discussed in Section˜E.2 (PN5) for $m_{1}\in\{10^{5},10^{6},10^{7}\}\,M_{\odot}$. All inspirals are evolved for 44 years, with $a=0.998$ and $\epsilon=10^{-5}$. The white dashed contours indicate regions where the dephasing crosses $\{0.01,0.1,1\}\,\mathrm{rad}$. The blue dashed line indicates the outer boundary in $(p_{0},e_{0})$ where the body plunges within four years in the PN5 model.
@@
Some of these PN5 results have already been integrated into few. For instance, the PN5-AAK waveform model combines the PN5 inspiral trajectory with the semi-relativisitic (“kludge”) amplitudes chua2017augmented as described in Ref. Katz:2021yft , providing a semi-relativistic waveform model defined over the generic inspiral parameter space: the few implementation of the PN5 mode amplitudes is also under development (cf. Section˜VI.2).
@@
    Validating the interpolations of the forcing functions $\hat{f}^{(0)}_{p,e}$ in the trajectory module against their PN5 equivalents (Fig.˜25).
@@
    Validating the interpolations of the mode amplitudes $\mathcal{A}_{\ell mn}$ against their PN5 equivalents for sample harmonic modes; the PN5 spherical harmonic waveform mode amplitudes were obtained from the PN5 Teukolsky spheroidal harmonic mode amplitudes via Eq.˜6 and projected onto spherical harmonics as described in Eq.˜11 (Fig.˜26).
@@
    Assessing the orbital dephasing (over four years) between the (fully-relativistic) trajectory model in few v2 and the PN5 inspiral model scales(Fig.˜27).
To do this, we have since implemented two minor improvements to the PN5 trajectory module. Firstly, we have corrected minor typos in the (adiabatic) PN5 fluxes of the angular momentum and Carter constant and the associated PN5 forcing functions $\hat{f}^{\mathrm{PN5}}_{p,e}$. Secondly, unlike the public database at BHPC BHPClub , $\hat{f}^{\mathrm{PN5}}_{p,e}$ here are computed from the PN5 fluxes of energy and angular momentum (in the equatorial limit) without expanding the Jacobian elements in Eq.˜7 in powers of $v$ or $e$, as doing so enhances the accuracy of $\hat{f}^{\mathrm{PN5}}_{p,e}$ in comparison to their numerical counterparts (cf. Ref. Isoyama:2021jjd ).
We find in Fig.˜25 that the agreement between the interpolations of the forcing functions $\hat{f}^{\mathrm{FEW}}_{p,e}$ and the PN5 forcing functions $\hat{f}^{\mathrm{PN5}}_{p,e}$ improves as $p$ increases and $e$ decreases. This agreement in ${\hat{f}}^{(0)}_{e}$ breaks down for small $e$, since both magnitudes of $\hat{f}^{\mathrm{PN5,Few}}_{e}$ become very small. A similar trend can be observed in the amplitude comparisons displayed in the top two panels of Fig.˜26, aside from the bottom left panel where the magnitude of the higher multipole amplitude varies rapidly and become small, and the bottom right panel where the accuracy of the PN5 amplitudes decrease closer to the separatrix. However, we find the interpolation error is predominant at smaller values of $e$ in this case. The deterioration of the agreement in ${\mathcal{A}}_{\ell mn}$ as the eccentricity becomes larger (typically $e\gtrsim 0.4$) is a result of the eccentricity expansion in the PN5 amplitudes, which become significantly less accurate in that region. This observation has been corroborated through a comparison of the PN5 amplitudes with numerical amplitude data from both the BHPT BHPToolkit and the BHPC Isoyama:2021jjd ; BHPClub . For weak-field trajectories with initial separation $p_{0}\gg p_{\text{sep}}$ in Fig.˜27, we observe that the dephasing between both models tends to zero (independent of the initial eccentricity $e_{0}$), indicating that the two models converge in the weak field limit, as expected.